OpenStudy (anonymous):

If I have the x-intercepts (-10,0) and (10,0) and the vertex (0,12), what would the equation of the parabola be in the form of y=A(x-h)^2+k?

OpenStudy (anonymous):

A determines if it opens up or down, in this case down.

OpenStudy (anonymous):

I know its down by plotting the vertex and x ints. this means A is less than 0

OpenStudy (anonymous):

But what is the equation? Where do the coordinates fit in?

OpenStudy (anonymous):

Is anyone going to help me?

OpenStudy (anonymous):

plot the point first

OpenStudy (anonymous):

a parabola is a "U" shape

OpenStudy (anonymous):

Isn't it inverted though?

OpenStudy (anonymous):

It opens down

OpenStudy (anonymous):

whats the answer choices I could solve it in 1 minute with them

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\color{blue}{\text{vertex(h,k)=(0,12),x intercepts (-10,0) and (10,0)}}\\~\\ \color{blue}{\text{u have to find the value of "a" here}}\\~\\ y=a(x-h)^2+k\\~\\ y=a(x-0)^2+12\\~\\ y=ax^2+12\\~\\ \color{blue}{\text{now put any one value of x intercept}}\\~\\ y=ax^2+12\\~\\ 0=a(-10)^2+12\\~\\ a=\dfrac{-12}{100}\\~\\ a=\dfrac{-3}{25}\\~\\ \color{blue}{\text{equation of parabola}}\\~\\ y=a(x-h)^2+k\\~\\ y=\dfrac{-3}{25}x^2+12 }\end{align}\)

OpenStudy (mathmath333):